# Multi-Agent Reinforcement Learning for Task Offloading in Wireless Edge Networks

### v1

@@ A.7 Proof of lemma 2
In order to prove the desired result we need to use theorem 2.
Due to the structure of the problem, we have two cases to prove; in particular, $t=\theta_{i}$ and $t=\theta_{j}$ for $j\neq i$. Note how we are studying the differentiability of $\tilde{J}_{i}$, i.e. the function related to agent $i$.
For both cases, consider the following:
  • $X=L^{\gamma}(\beta)\times[0,1]^{N-1}$: this is the product between the set of stationary distributions and the possible values of the constraints (excluding $t$). Its elements will be denoted as $(\rho,\theta_{-j})\in L^{\gamma}(\beta)\times[0,1]^{N-1}$ assuming that $t=\theta_{j}$
@@
  • $f(x,t)=f(\rho,\theta_{-i},\theta_{i})=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i})$, with $r_{i}:\mathcal{S}\times\mathcal{A}\times[0,1]^{N-1}\to\mathbb{R}$ continuous wrt $\theta_{-i}$
  • $g(x,t)=g(\rho,\theta_{-i},\theta_{i})=\sum_{s,a}\rho(s,a)c(s,a)-\frac{\theta_{i}}{\theta_{max}}$, with with $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{N}$
@@
    $L(\rho,\lambda,\theta)=f(\rho,\theta)+\lambda_{i}g(\rho,\theta)$
@@
$f$ and $g$ are continuous and concave in $x$
@@
$f_{t}(x,t)$ and $g_{t}(x,t)$ are continuous in $(x,t)$
  • $t=\theta_{i}$: the immediate reward function does not depend on $\theta_{i}$, therefore $f_{t}(\cdot)=0$. On the other hand, $g_{t}(\cdot)=\frac{1}{\theta_{max}}$
  • $t=\theta_{j},\ j\neq i$: the continuity of $f_{t}(x,t)$ in $(x,t)$ is an immediate consequence of the definition of the reward function $r_{i}$ and the assumption that it is continuous with regard to $\theta_{-i}$ and therefore it is differential wrt $\theta_{j},\ \forall j\neq i$. On the other hand, $g_{t}(\cdot)=0$
Existence of $\hat{x}\in X$ such that $g(\hat{x},t)\gg 0$ for all $t\in[0,1]$
This allows us to conclude that the function $\tilde{J}_{i}(\theta)$ is differentiable wrt every component of $\theta$ and therefore that theorem 1 is verified for the problem studied.
Finally, note how we can use Theorem 3.3 in [3] to show how an optimal solution $\rho^{\star}$ for LP is such that the stationary policy $\pi(\rho^{\star})$ is optimal for the original constrained problem. ∎
A.8 Proof of proposition 2
@@
In our context, for agent $i$, the function $V(t)$ corresponds to $\hat{J}_{i}(\theta)$, while the function $L$ represents the Lagrangian for agent $i$
$\mathcal{L}_{i}(\rho,\lambda,\theta_{i},\theta_{-i})=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i})+\lambda\left(\sum_{s,a}\rho_{i}(s,a)c_{i}(s,a)-\frac{\theta_{i}}{\theta_{max}}\right)$
When considering the parameter $t=\theta_{i}$, it is easily verified that $\frac{\partial\mathcal{L}_{i}}{\partial\theta_{i}}=-\frac{\lambda}{\theta_{max}}$ and therefore when $\theta_{i}\in(0,\theta_{max})$ it is verified
@@
On the other hand, when $t=\theta_{j}$, with $j\neq i$
$\frac{\partial}{\partial\theta_{j}}\mathcal{L}_{i}(\rho,\lambda,\theta_{i},\theta_{-i})$ $=\frac{\partial}{\partial\theta_{j}}\sum_{s_{i},a_{i}}\rho_{i}(s_{i},a_{i})\hat{r}_{i}(s_{i},a_{i})$
@@
$\frac{\partial}{\partial\theta_{j}}\tilde{J_{i}^{\ell}}(\theta_{i},\theta_{-i})=$ $\max_{\rho}\min_{\lambda}\sum_{s_{i}}\rho(s_{i},a_{crowd})\frac{\partial}{\partial\theta_{j}}d(1+\theta_{-i})$
@@ C.3 Comparison with MAPPO over longer timeframes
C.4 Effect of Initial Constraint
C.5 Additional algorithms for fig. 1(a)
@@ C.4 Effect of Initial Constraint
C.6 Non linear penalty
C.7 Evaluation gradient objective function
C.8 Verification of the Approximate Reward Bound
@@ Appendix D Extended Algorithm
3:  for $k=0,1,\dots,K(n)$ do $\triangleright$ Intermediate timescale — $\lambda$ update
4:   for $t=0,1,\dots,T(k,n)$ do $\triangleright$ Fastest timescale — policy update
5:     Collect trajectories using policy $\pi_{i}^{t,k,n}$
@@
8:     Update policy: $\pi_{i}^{t+1,k,n}\leftarrow$ RL step on $r_{i}^{\text{total}}$
@@
10:   Estimate constraint $K_{i}(\pi_{i}^{\cdot,k,n})$
11:   $\lambda_{i}^{k+1,n}\leftarrow\lambda_{i}^{k,n}+\alpha_{k}(K_{i}-\theta_{i})$
@@
13:  return $\tilde{J}^{\ell}_{i}(\theta_{i},\theta_{-i})\leftarrow$ evaluate local policy $\pi_{i}^{T(K(n),n),K(n)}$
@@
19:   $\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta^{n}_{-i})\leftarrow$ evaluate_optimal_constrained_policy($i,\theta_{i}^{n},\theta_{-i}^{n},n$)
@@
23:   $\tilde{J}^{\ell}_{i}(\theta_{i}^{n}+\epsilon_{i},\theta^{n}_{-i})\leftarrowevaluate$_optimal_constrained_policy($i,\theta_{i}^{n}+\epsilon_{i},\theta_{-i},n$)
24:   $\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta^{n}_{-i}+\epsilon_{i})\leftarrowevaluate$_optimal_constrained_policy($i,\theta_{i}^{n},\theta_{-i}+\epsilon_{i},n$)
@@
27:   $\frac{\partial}{\partial\theta_{i}}\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\leftarrow\left(\tilde{J}^{\ell}_{i}(\theta_{i}^{n}+\epsilon_{i},\theta_{-i}^{n})-\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\right)/\epsilon_{i}$
28:   $\frac{\partial}{\partial\theta_{-i}}\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\leftarrow\left(\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta_{-i}^{n}+\epsilon_{i})-\tilde{J}^{\ell}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\right)/\epsilon_{i}$
29:   $\nabla\frac{\partial}{\partial\theta_{i}}\tilde{J}^{\ell}(\theta)=\frac{\partial\tilde{J}^{\ell}_{i}(\theta)}{\partial\theta_{i}}+\sum_{j\neq i}\frac{\partial\ \tilde{J}^{\ell}_{j}(\theta)}{\partial\theta_{-j}}$
30:   $\theta_{i}^{n+1}\leftarrow\theta_{i}^{n}+\eta_{n}\nabla_{\theta_{i}}\tilde{J}^{\ell}_{i}(\theta_{i}^{n})$
@@
Appendix E Numerical experiments hyperparameters

### v2

@@ A.3 Proof of lemma 2
In order to prove the desired result for the function $\hat{J}_{i}(\theta)$ we need to use theorem 2 and make a distinction between two separate cases, $t=\theta_{i}$ and $t=\theta_{j}$.
We consider
$V(t)=\min_{y\in\mathbb{R}^{k}_{+}}\max_{x\in X}L(x,y,t)=\max_{x\in X}\min_{y\in\mathbb{R}^{k}_{+}}L(x,y,t)$
Next we define how all the quantities involved in the two different cases.
A.3.1 $\boldsymbol{t=\theta_{i}}$
In this case, all the values of $\theta_{j}$ for $j\neq i$ are fixed and simply define the function $f$.
Consider the following:
  • $X=L^{\gamma}(\beta)$: this is the set of stationary distributions given an initial distribuiton $\beta$ and the discount factor $\gamma$
@@ A.3.1 $\boldsymbol{t=\theta_{i}}$
  • $f(x,t)=f_{\theta_{-i}}(\rho)=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i})$, with $\hat{r}_{i}:\mathcal{S}\times\mathcal{A}\times[0,1]^{N-1}\to\mathbb{R}$ defined by the fixed parameters $\theta_{-i}$
  • $g(x,t)=g(\rho,\theta_{i})=\sum_{s,a}\rho(s,a)c(s,a)-\frac{\theta_{i}}{\theta_{max}}$, with with $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{N}$
@@
    $L_{i}(\rho,\lambda,\theta_{i})=f_{\theta_{-i}}(\rho)+\lambda_{i}g(\rho,\theta_{i})$
    assuming we are studying the differentiability wrt $\theta_{i}$
Now we show that the hypotheses of Theorem 2 are verified.
X is a convex compact set in a normed linear space
$\boldsymbol{f}$ and $\boldsymbol{g}$ are continuous and concave in $\boldsymbol{x}$
$\boldsymbol{f_{t}(x,t)}$ and $\boldsymbol{g_{t}(x,t)}$ are continuous in $\boldsymbol{(x,t)}$
The immediate reward function does not depend on $\theta_{i}$, therefore $f_{t}(\cdot)=0$. On the other hand, $g_{t}(\cdot)=\frac{1}{\theta_{max}}$
Existence of $\boldsymbol{\hat{x}\in X}$ such that $\boldsymbol{g(\hat{x},t)\gg 0}$ for all $\boldsymbol{t\in[0,1]}$
A.3.2 $\boldsymbol{t=\theta_{j}}$
In this case, all the values of $\theta_{k}$ for $k\neq j$ are fixed; they will be considered fixed parameters in the definition of $f$ and $g$. Consider the following:
  • $X=L^{\gamma}(\beta)$ this is the set of stationary distributions
  • $f(x,t)=f_{\theta_{-i,j}}(\rho,\theta_{j})=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i,j},\theta_{j})$, with $\theta_{-i,j}=\sum_{k\neq i,j}\theta_{k}$
  • $g(x,t)=g_{\theta_{i}}(\rho)=\sum_{s,a}\rho(s,a)c(s,a)-\frac{\theta_{i}}{\theta_{max}}$, with with $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{N}$
  • the Lagrangian is explicitly written as
    $L_{i}(\rho,\lambda,\theta_{j})=f_{\theta_{-i,j}}(\rho,\theta_{j})+\lambda_{i}g_{\theta_{i}}(\rho)$
@@ A.3.2 $\boldsymbol{t=\theta_{j}}$
$\boldsymbol{f}$ and $\boldsymbol{g}$ are continuous and concave in $\boldsymbol{x}$
@@
$\boldsymbol{f_{t}(x,t)}$ and $\boldsymbol{g_{t}(x,t)}$ are continuous in $\boldsymbol{(x,t)}$
The continuity of $f_{t}(x,t)$ in $(x,t)$ is an immediate consequence of the definition of the reward function $r_{i}$ and the assumption that it is continuous with regard to $\theta_{-i}$ and therefore it is differential wrt $\theta_{j},\ \forall j\neq i$. On the other hand, $g_{t}(\cdot)=0$
Existence of $\boldsymbol{\hat{x}\in X}$ such that $\boldsymbol{g(\hat{x},t)\gg 0}$ for all $\boldsymbol{t\in[0,1]}$
The existence of a point $\hat{x}\in X$ such that $g(\hat{x},t)\gg 0$ for all $t\in[0,1]$ corresponds to the existence of a stationary distribution that strictly satisfies all constraints. This condition is met, for example, by a policy that never selects the crowded action—ensuring all constraints are strictly satisfied—provided that $\theta_{j}>0$.
A.3.3 Conclusion of the proof
Thanks to what is showed in sections A.3.1 and A.3.2 we know that the objective function $V(t)$ is absolutely continuous with regard to $t$, for all choices of $t$.
Consider now the function
$\hat{J}_{i}^{\rho}(\theta)=\mathbb{E}_{s,a\sim\rho^{\star}_{i}}\left[\hat{r}_{i}(s,a;\theta_{-i)}\right]$
where $\rho^{\star}_{i}$ depends on $\theta$.
When we fix all coordinates $\theta_{k}$ for $k\neq j$, the mapping
$\theta_{j}\mapsto\hat{J}_{i}^{\rho}(\theta)$
coincides with $V(\theta_{j})$, which is known to be absolutely continuous.
It then follows that $\hat{J}_{i}^{\rho}(\theta)$ is differentiable a.e.
Finally, note how we can use Theorem 3.3 in [3] to show how an optimal solution $\rho^{\star}$ for LP is such that the stationary policy $\pi(\rho^{\star})$ is optimal for the original constrained problem. This implies that, from the optimal stationary distribution we can also retrieve the optimal policy.
This allows us to conclude that the function $\hat{J}_{i}(\theta)$ is differentiable wrt every component of $\theta$ and therefore that theorem 1 is verified for the problem studied.
A.4 Proof of proposition 3
@@
When considering the parameter $t=\theta_{i}$ and $\theta_{j}$ fixed for $j\neq i$, we know that
$L(\rho,\lambda,\theta_{i})$ $=f_{\theta_{-i}}(\rho)+\lambda_{i}g(\rho,\theta_{i})$
$=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i,j},\theta_{j})+\lambda_{i}\left(\sum_{s,a}\rho(s,a)c(s,a)-\frac{\theta_{i}}{\theta_{max}}\right)$
It is easily verified that $\frac{\partial L(\theta_{i})}{\partial\theta_{i}}=-\frac{\lambda}{\theta_{max}}$ and therefore when $\theta_{i}\in(0,\theta_{max})$ it is verified
@@
On the other hand, when $t=\theta_{j}$, with $j\neq i$, we have
$L(\rho,\lambda,\theta_{j})$ $=f_{\theta_{-i,j}}(\rho,\theta_{j})+\lambda_{i}g_{\theta_{i}}(\rho)$
$=\sum_{s,a}\rho_{i}(s,a)\hat{r}_{i}(s,a;\theta_{-i,j},\theta_{j})+\lambda_{i}\left(\sum_{s,a}\rho(s,a)c(s,a)-\frac{\theta_{i}}{\theta_{max}}\right)$
It then follows that
$\frac{\partial}{\partial\theta_{j}}L(\rho,\lambda,\theta_{j})$ $=\frac{\partial}{\partial\theta_{j}}\sum_{s_{i},a_{i}}\rho_{i}(s_{i},a_{i})\hat{r}_{i}(s_{i},a_{i})$
@@
$\frac{\partial}{\partial\theta_{j}}\hat{V}(\theta_{j})=$ $\max_{\rho}\min_{\lambda}\sum_{s_{i}}\rho(s_{i},a_{crowd})\frac{\partial}{\partial\theta_{j}}d(1+\theta_{-i})$
@@
where the optimal stationary distribution $\rho^{\star}$ is a function of the constraint vector $\theta$. ∎
@@ C.3 Comparison with MAPPO over longer timeframes
C.4 Additional algorithms for fig. 1(a)
@@
C.5 Verification of the Approximate Reward Bound
@@ Appendix D Extended Algorithm
The following pseudocode corresponds to the version of the algorithm that uses finite differences to compute the gradient of the objectove function. This is the version used in the nu,erical experiments.
@@
3:  for $m=0,1,\dots,M(n)$ do $\triangleright$ Intermediate timescale — $\lambda$ update
4:   for $t=0,1,\dots,T(m,n)$ do $\triangleright$ Fastest timescale — policy update
5:     Collect trajectories using policy $\pi_{i}^{t,m,n}$
@@
8:     Update policy: $\pi_{i}^{t+1,m,n}\leftarrow$ RL step on $r_{i}^{\text{total}}$
@@
10:   $\hat{K}_{i}\leftarrow$ $K_{i}(\pi_{i}^{\cdot,m,n})$ $\triangleright$ Estimate long term cost of current policy
11:   $\lambda_{i}^{m+1,n}\leftarrow\lambda_{i}^{m,n}+\alpha_{k}(\hat{K}_{i}-\theta_{i})$
@@
13:  return $\hat{J}_{i}(\theta_{i},\theta_{-i})\leftarrow$ evaluate local policy $\pi_{i}^{T(M(n),n),M(n)}$
@@
19:   $\hat{J}_{i}(\theta_{i}^{n},\theta^{n}_{-i})\leftarrow$ evaluate_optimal_constrained_policy($i,\theta_{i}^{n},\theta_{-i}^{n},n$)
@@
23:   $\hat{J}_{i}(\theta_{i}^{n}+\epsilon_{i},\theta^{n}_{-i})\leftarrowevaluate$_optimal_constrained_policy($i,\theta_{i}^{n}+\epsilon_{i},\theta_{-i},n$)
24:   $\hat{J}(\theta_{i}^{n},\theta^{n}_{-i}+\epsilon_{i})\leftarrowevaluate$_optimal_constrained_policy($i,\theta_{i}^{n},\theta_{-i}+\epsilon_{i},n$)
@@
27:   $\frac{\partial}{\partial\theta_{i}}\hat{J}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\leftarrow\left(\hat{J}_{i}(\theta_{i}^{n}+\epsilon_{i},\theta_{-i}^{n})-\hat{J}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\right)/\epsilon_{i}$
28:   $\frac{\partial}{\partial\theta_{-i}}\hat{J}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\leftarrow\left(\hat{J}_{i}(\theta_{i}^{n},\theta_{-i}^{n}+\epsilon_{i})-\hat{J}_{i}(\theta_{i}^{n},\theta_{-i}^{n})\right)/\epsilon_{i}$
29:   $\nabla\frac{\partial}{\partial\theta_{i}}\hat{J}(\theta)=\frac{\partial\hat{J}_{i}(\theta)}{\partial\theta_{i}}+\sum_{j\neq i}\frac{\partial\ \hat{J}_{j}(\theta)}{\partial\theta_{-j}}$
30:   $\theta_{i}^{n+1}\leftarrow\theta_{i}^{n}+\eta_{n}\nabla_{\theta_{i}}\hat{J}_{i}(\theta_{i}^{n})$
@@
Appendix E Extended Algorithm with exact evaluations of $\boldsymbol{\lambda_{i}^{\star}}$
The following pseudocode corresponds to the version of the algorithm that assumes access to the exact values of $\lambda_{i}^{\star}$, obtained through the learning process, for computing the gradient.
Algorithm 3 DCC Framework
2:function OptimizeLocalCMDP($i,\theta,n$)
3:  for $m=0,1,\dots,M(n)$ do $\triangleright$ Intermediate timescale — $\lambda$ update
4:   for $t=0,1,\dots,T(m,n)$ do $\triangleright$ Fastest timescale — policy update
5:     Collect trajectories using policy $\pi_{i}^{t,m,n}$
7:
8:     Update policy: $\pi_{i}^{t+1,m,n}\leftarrow$ RL step on $r_{i}^{\text{total}}$
9:   end for
10:   $\hat{K}_{i}\leftarrow$ $K_{i}(\pi_{i}^{\cdot,m,n})$ $\triangleright$ Estimate long term cost of current policy
11:   $\lambda_{i}^{m+1,n}\leftarrow\lambda_{i}^{m,n}+\alpha_{k}(\hat{K}_{i}-\theta_{i})$
12:  end for
13:  $\hat{J}_{i}(\theta)\leftarrow$ evaluate long term reward of policy $\pi_{i}^{T(M(n),n),M(n)}$
14:  return $\hat{J}_{i}(\theta),\lambda_{i}^{M(n),n},\pi_{i}^{T(M(n),n),M(n),n}$
15:end function
16:
17:
18:Compute $\theta_{i}^{MAX}$ for each agent $i$: $\theta_{i}^{MAX}\leftarrow K_{i}(\pi_{i}^{\text{always offload}})$
19:for $n=0,1,2,\dots$ do $\triangleright$ Slowest timescale — constraint update
20:  for each agent $i=1,\dots,N$ do
21:   $\hat{J}_{i}(\theta^{n}),\lambda_{i}^{\star},\pi_{i}^{\star}\leftarrowOptimizeLocalCMDP$($i,\theta^{n},n$)
22:  end for
23:  for each agent $i$ do
24:   $\frac{\partial}{\partial\theta_{i}}\hat{J}_{i}(\theta^{n})\leftarrow-\frac{\lambda_{i}^{\star}}{\theta_{i}^{MAX}}$
25:   for each agent $j\neq i$ do
26:     $\frac{\partial}{\partial\theta_{j}}\hat{J}_{i}(\theta^{n})\leftarrow\theta_{i}\frac{\partial}{\partial\theta_{j}}d(1+\theta_{-i})$
27:   end for
28:  end for
29:  for each agent $i=1,\dots,N$ do
30:   $\nabla\frac{\partial}{\partial\theta_{i}}\hat{J}(\theta)=\frac{\partial\hat{J}_{i}(\theta)}{\partial\theta_{i}}+\sum_{j\neq i}\frac{\partial\ \hat{J}_{j}(\theta)}{\partial\theta_{i}}$
31:   $\theta_{i}^{n+1}\leftarrow\theta_{i}^{n}+\eta_{n}\nabla_{\theta_{i}}\hat{J}_{i}(\theta_{i}^{n})$
32:  end for
33:end for
Appendix F Numerical experiments hyperparameters